OpenStudy (anonymous):

How do I show a series is convergent and what it converges to? The series is \[1 + \frac{x^n}{n!} ...\]

OpenStudy (jamesj):

You mean 1 + x + x^2/2!d + x^3/3! + x^4/4! + ... ?

OpenStudy (anonymous):

Yea

OpenStudy (jamesj):

You mean 1 + x + x^2/2! + x^3/3! + x^4/4! + ... ?

OpenStudy (jamesj):

There are lots of convergence tests. Which ones do you know so far?

OpenStudy (anonymous):

I don't know any. We're supposed to figure it out on our own but I don't really understand

OpenStudy (jamesj):

Well, when you have a geometric progression, when does this converge: a + ar + ar^2 + ar^3 + ar^4 + ar^5 + ... ?

OpenStudy (anonymous):

What are 'a' and 'r'? Both variables?

OpenStudy (jamesj):

Yes, you recognize this as a geometric series, yes? Such as 1 + 1/2 + 1/4 + 1/8 + 1/16 + ... in which case a = 1 and r = 1/2

OpenStudy (anonymous):

Oh, yea. Convergence just means that the entire series can be represented by a number, right?

OpenStudy (jamesj):

Yes, it means 1 + 1/2 + 1/4 + 1/8 + 1/16 + ... gets arbitrarily close to a given number if we have enough terms of the sum. In this case, the sum converges to 2.

OpenStudy (anonymous):

I have that memorized but I forgot why it's true or the proof for it

OpenStudy (jamesj):

Given that you are having trouble with a geometric series, I have a hard time that you are being asked to come up with your own convergence test. Are you sure you haven't missed a class/lecture which you need to look up in your text book?

OpenStudy (jamesj):

...have a hard time believing that ...

OpenStudy (anonymous):

Our last lesson was on inverse functions. Our teacher makes us learn the lesson before he actually teaches it. I don't think we cover series until the very end of the semester because we have to go over more functions and stuff

OpenStudy (jamesj):

Well, the proof is instructive: The finite sum a + ar + ar^2 + .... + ar^n is equal to a (1 - r^(n+1)) / (1 - r) Look familiar?

OpenStudy (anonymous):

I've seen it before

OpenStudy (jamesj):

Now what is the limit of r^(n+1) as n --> infinity ?

OpenStudy (anonymous):

Doesn't it depend on whether r is <1, =1, or >1

OpenStudy (jamesj):

Yes: if |r| < 1, |r| = 1 or |r| > 1. For what values of r does the limit r^(n+1) exist as n --> infinity?

OpenStudy (anonymous):

I think all real numbers

OpenStudy (anonymous):

Oh wait

OpenStudy (jamesj):

No. If r = 1, for example with a = 1, we have 1 + 1 + 1 + 1 .... which does not converge. In other words \[ S_n = a \frac{1-r^{n+1}}{1-r} \] Does not converge when n --> infinity

OpenStudy (anonymous):

i don't think it has a limit if the number fluctuates

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Why doesn't limit 1^(n+1) as n approaches infinity = 1?

OpenStudy (jamesj):

In fact, S_n as we're written it doesn't even exist when r = 1 because the denominator is zero. But yes indeed, \[ \lim_{n \rightarrow \infty} r^{n+1} \] exists provided \( -1 < r \leq 1 \) But now we just saw that if r = 1, this is a special case and it is clear that if a is not zero, then a + ar + ar^2 + ar^3 + .... doesn't exist. Hence in summary...

OpenStudy (jamesj):

The infinite sum a + ar + ar^2 + ar^3 + ... exists if and only if |r| < 1 Yes?

OpenStudy (anonymous):

Yes

OpenStudy (jamesj):

Now, label each of these terms like this \[ a_0 = a \] \[ a_1 = ar \] \[ a_2 = ar^2 \] ... \[ a_j = ar^j \] ... Then it is clear that for all \( j > 0 \), \( a_j/a_{j-1} = r \)

OpenStudy (jamesj):

Therefore by comparing a sum with a geometric sum, \[ a_0 + a_1 + a_2 + .... + a_j + .... \] converges if and only if \[ \lim_{n \rightarrow \infty} \frac{a_n}{a_{n-1}} < 1 \]

OpenStudy (jamesj):

In the case of the series you wrote down \[ a_0 = 1, a_1 = x, a_2 = x^2/2!, ..., a_j = x^j/j!, ... \]

OpenStudy (jamesj):

Therefore \[ 1 + x + x^2/2! + .... = a_0 + a_1 + a_2 + ... \] converges if and only if \[ \lim_{n \rightarrow \infty} \frac{a_n}{a_{n-1}} < 1 \] i.e., \[ \lim_{n \rightarrow \infty} \frac{x^n/n!}{x^{n-1}/(n-1)!} \ = \ \lim_{n \rightarrow \infty} \frac{x}{n} \ < \ 1 \]

OpenStudy (jamesj):

As this is true for all real numbers x, it must be that 1 + x + x^2/2! + x^3/3! + ... converges for all real numbers x

OpenStudy (anonymous):

Wow, that is pretty cool

OpenStudy (jamesj):

...and not at all obvious if you've not seen it before. I'm surprised you're asked to figure this out.

OpenStudy (jamesj):

Now the second part: what is this equal to? Call this function f(x) f(x) = 1 + x + x^2/2! + x^3/3! + ... Now what is df/dx ?

OpenStudy (anonymous):

df/dx is 0 + 1 + x + x^2/2 + x^3/6, etc. or the same thing as f(x) I guess? since 6 = 3! I kind of see where this is going

OpenStudy (jamesj):

Yes, f' = f What function behaves like this?

OpenStudy (anonymous):

e

OpenStudy (jamesj):

e^x

OpenStudy (anonymous):

Oh right

OpenStudy (anonymous):

But why did you differentiate to find out what the value it converged to was?

OpenStudy (jamesj):

There's one little twist. So also does g(x) = Ae^x for any constant A. How do we get A = 1?

OpenStudy (jamesj):

Why? Because it's the easiest method right now. You'll see later how to start with a function and arrive at a function.